OpenStudy (anonymous):

Help me now! PLEASE Solve the following simultaneous equation using the substitution method: y+1=2x y=x+2 ASAP please!

OpenStudy (anonymous):

y+1=2x replace y with x+2: x+2 + 1 = 2x x=3 y=2x-1 y=5

OpenStudy (anonymous):

well that the answer :) good luck

OpenStudy (anonymous):

Help me with this one as well! 3x+8y=24 x+y=3 pleaseeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

ANyone?

OpenStudy (anonymous):

i only know that one by elimination sorry

OpenStudy (anonymous):

oh sorry i thought that it was a plus sign now thats possible

OpenStudy (anonymous):

Help me? please? and quick?:S

OpenStudy (anonymous):

take the second equation an make x the subject of the formula x=3/y Afterwards replace the x value with 3/y to find y. After finding y raka u then substitute y value on the secon equation to find x simple:)

OpenStudy (anonymous):

Can you just do the whole just like Abermejo did on top? :S please?

OpenStudy (anonymous):

x+y=3 tomake x the subject we have x=3-y Now since 3x/8y=24 =3(3/y/8y)=24 =9/y/8y=24 simplying this further we get 9/y*8y=24 =8y^2/9=24 and then multiply by 9 both sides tog 8y^2=216 the divide by 8 both sides to get y^2=27 then take the sqrt both sides to get your y as=5.196 =5.2 Afterward u take this y value put it in the second equation to find ur x value x+y=3 x+5.2=3 x=-2.2

OpenStudy (anonymous):

please copy this down in ur paper for u to understand properly

OpenStudy (anonymous):

feel free to ask any unclear thing

OpenStudy (anonymous):

rika a u still there.sry I made a huge blander i thought that it was + was division sign again. y=3-x from second equation now take that y and replace it with the y on the first equation =3x+3-x=24 =2x=21 x=21/2 x=10.5 for your y x+y=3 10.5+y=3 y=7.5

OpenStudy (anonymous):

y=-7.5 sry

OpenStudy (anonymous):

answer is apparently wrong :\ y is 3 and x is 0

OpenStudy (anonymous):

there are many ways of killing a cat substitute my x and y valu on the second equation u`ll still get 3 10+(-7.5)=3

OpenStudy (anonymous):

Alright thanks pal

OpenStudy (anonymous):

Gud luck :)

OpenStudy (anonymous):

:)